OpenStudy (anonymous):

5^5x-3 = 31 What is x=?

OpenStudy (mertsj):

\[(5x-3)\log5 = \log 31\]

OpenStudy (anonymous):

hmmmm?

OpenStudy (mertsj):

\[(5x-3)(.699)=1.4914\]

OpenStudy (mertsj):

\[3.495x-2.097=1.4914\]

OpenStudy (mertsj):

You can finish it.